OpenStudy (anonymous):

One number is eight less than another number. The sum of the numbers is eighty. Find both numbers.

OpenStudy (anonymous):

plz help me figure this out

OpenStudy (anonymous):

Set up your problem just like it sounds, One number is 8 less than another number. 'N' being the original number, and 'N-8' being the second (But, in the problem I set up it will be called A until I substitute). their sum is 80. So your problem looks like N + (A) = 80 A = N- 8 So you plug in... N + (N-8) = 80... Solve from there. 2N - 8 = 80 2N = 88 N = 44 A = (44) - 8 A = 36 Check: 44+ 36 = 80.

OpenStudy (anonymous):

Thank you so much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!